# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ D.1 Proof of Theorem 5: concentration inequalities of the M-steps at the true parameter
Taking the union bounds over $\ell\in[L]$, choosing $t=\log(n)$ and using $n\pi_{\min}^{*}\geq C\log n$ yield (4.2).
@@
Recalling the model (1.1), we partition $n]=\bigcup_{a=1}^{L}\widehat{G}_{a}$ with
Further let
Pick any $\ell\in[L]$. Notice that
We distinguish two cases depending on the order of $\Delta_{\min}$.
Case 1: $\Delta_{\min}\leq C\log n$.
Further observing that
$\sum_{b=1}^{L}\pi_{b}^{*}\mathbb{E}[V_{\ell,i}\mid Y_{i}=b]=\sum_{b=1}^{L}\pi_{b}^{*}~\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})(X_{i}-\mu_{\ell}^{*})\mid Y_{i}=b\right]=\mathbb{E}[\gamma_{\ell}(X_{i};\omega^{*})](\bm{\mu}_{\ell}(\omega^{*})-\mu_{\ell}^{*})=0$
adding and subtracting terms gives $\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}\leq\textrm{I}_{\ell}+\textrm{II}_{\ell}$ where
$\textrm{I}_{\ell}$ $=~{1\over n\widehat{\bm{\pi}}_{\ell}(\omega^{*})}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\left(V_{\ell,i}-\mathbb{E}[V_{\ell,i}\mid Y=a]\right)\right\|_{\Sigma^{*}}$
$\textrm{II}_{\ell}$ $={1\over\widehat{\bm{\pi}}_{\ell}(\omega^{*})}\left\|{\sum_{a=1}^{L}\left({n_{a}\over n}-\pi_{a}^{*}\right)\mathbb{E}[V_{\ell,i}\mid Y_{i}=a]}\right\|_{\Sigma^{*}}.$
We proceed to bound each term on the event
@@
To bound $\textrm{II}_{\ell}$, it is easy to see that
${1\over\widehat{\bm{\pi}}_{\ell}(\omega^{*})}\left\|\left(\widetilde{\pi}_{\ell}-\pi_{\ell}^{*}\right)\mathbb{E}[V_{\ell,i}\mid Y_{i}=\ell]\right\|_{\Sigma^{*}}$ $\overset{\mathcal{E}_{\pi}}{\lesssim}\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega^{*})N\mid Y=\ell\right]\right\|_{2}\sqrt{\log n\over n\pi_{\ell}^{*}}$
$\leq\sup_{v\in\mathbb{S}^{d}}\sqrt{\mathbb{E}[(v^{\top}N)^{2}]}\sqrt{\log n\over n\pi_{\ell}^{*}}$
$=\sqrt{\log n\over n\pi_{\ell}^{*}}.$ (D.7)
For any $a\neq\ell$, since
$\|\mathbb{E}[\gamma_{\ell}(X;\omega^{*})(X-\mu_{\ell}^{*})\mid Y=a]\|_{\Sigma^{*}}$
$\leq~\|\mathbb{E}[\gamma_{\ell}(X;\omega^{*})N\mid Y=a]\|_{\Sigma^{*}}+\sqrt{\Delta_{a\ell}}\mathbb{E}[\gamma_{\ell}(X;\omega^{*})\mid Y=a]$
@@
we find that
${1\over\widehat{\bm{\pi}}_{\ell}(\omega^{*})}\left\|{\sum_{a\in[L]\setminus\{\ell\}}\left(\widetilde{\pi}_{a}-\pi_{a}^{*}\right)\mathbb{E}[V_{\ell,i}\mid Y_{i}=a]}\right\|_{\Sigma^{*}}$
$\leq\max_{a\in[L]}{|\widetilde{\pi}_{a}-\pi_{a}^{*}|\over\pi_{a}^{*}}{1\over\pi_{\ell}^{*}}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}^{*}\|\mathbb{E}[\gamma_{\ell}(X;\omega^{*})(X-\mu_{\ell}^{*})\mid Y=a]\|_{\Sigma^{*}}$
$\lesssim\sqrt{\log n\over n\pi_{\min}^{*}}\left(L+{1\over\pi_{\ell}^{*}}\right)\sqrt{\Delta_{\max}}\exp(-c\Delta_{\min}).$ (D.8)
Combining (D.1.2) and (D.8) and using (3.4) yield
$\textrm{II}_{\ell}\lesssim\sqrt{\log n\over n\pi_{\ell}^{*}}.$ (D.9)
We proceed to bound $\textrm{I}_{\ell}$ via a truncation argument together with the vector-valued Bernstein inequality. Consider the event
$\mathcal{E}_{2}=\bigcap_{i=1}^{n}\mathcal{E}_{2,i},\qquad\mathcal{E}_{2,i}:=\left\{\|N_{i}\|_{2}\leq\sqrt{d}+2\sqrt{\log n}\right\}$ (D.10)
and by Lemma 31, one has $\mathbb{P}(\mathcal{E}_{2,i})\geq 1-n^{-2}$ and $\mathbb{P}(\mathcal{E}_{2})\geq 1-n^{-1}$. It is easy to see that
${1\over n}\left\|\sum_{a\in[L]}\sum_{i\in\widehat{G}_{a}}\mathbb{E}[V_{\ell,i}1\{\mathcal{E}_{2,i}^{c}\}]\right\|_{\Sigma^{*}}\leq~\max_{i\in[n]}\mathbb{E}\left[\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}1\{\mathcal{E}_{2,i}^{c}\}\right]\lesssim\sqrt{d+\Delta_{\max}\over n^{2}}.$ (D.11)
@@
$\textrm{I}_{\ell,1}:={1\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\Bigl{\{}V_{\ell,i}1\{\mathcal{E}_{2,i}\}-\mathbb{E}\left[V_{\ell,i}1\{\mathcal{E}_{2,i}\}\mid Y_{i}=a\right]\Bigr{\}}\right\|_{\Sigma^{*}}.$
@@
$\max_{i\in[n]}\|V_{\ell,i}1\{\mathcal{E}_{2,i}\}\|_{\Sigma^{*}}\leq\sqrt{d}+2\sqrt{\log n}+\sqrt{\Delta_{\max}}.$
Since the variance term $\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}[\left\|V_{\ell,i}1\{\mathcal{E}_{2,i}\}-\mathbb{E}[V_{\ell,i}1\{\mathcal{E}_{2,i}\}\mid Y=a]\right\|_{\Sigma^{*}}^{2}]$ is no greater than
$\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})^{2}\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}1\{\mathcal{E}_{2,i}\}\right]$
$\leq\sum_{i\in\widehat{G}_{\ell}}\mathbb{E}\left[\|N_{i}\|_{2}^{2}\right]+\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}2\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})(\|N_{i}\|_{2}^{2}+\Delta_{a\ell})\right]$
$\lesssim dn_{\ell}+\sum_{a\in[L]\setminus\{\ell\}}n_{a}{\pi_{\ell}^{*}+\pi_{a}^{*}\over\pi_{a}^{*}}(d+\Delta_{a\ell})\exp{(-c\Delta_{\min})}$ by Lemma 22
$\lesssim dn\pi_{\ell}^{*}+n(L\pi_{\ell}^{*}+1)(d+\Delta_{\max})\exp{(-c\Delta_{\min})}$ $\text{by }\mathcal{E}_{\pi}$
@@
applying the vector-valued Bernstein’s inequality in Lemma 34, conditioning on $Y_{1},\ldots,Y_{n}$, gives that for all $t\geq 0$,
$\textrm{I}_{\ell,1}\lesssim\sqrt{t\over n}\sqrt{d\pi_{\ell}^{*}}+{t\over n}\sqrt{d+\log n+\Delta_{\max}}$
with probability at least $1-e^{-t}-2n^{-1}$. By the independence of $Y_{i}$ and $N_{i}$, the above also holds unconditionally. Together with (D.11) and (D.9), by choosing $t=\log n$ and using $n\pi_{\min}^{*}\geq C\log(n)(1\vee\log(n)/d)$, we obtain that for any $t\geq 1$, with probability $1-3n^{-1}$,
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}$ $\lesssim\sqrt{d\log n\over n\pi_{\ell}^{*}}+{\log n\over n\pi_{\ell}^{*}}\sqrt{\Delta_{\max}}\lesssim\sqrt{d\log n\over n\pi_{\ell}^{*}}+{\log^{3/2}n\over n\pi_{\ell}^{*}}\sqrt{\Delta_{\max}\over\Delta_{\min}}.$ (D.12)
The last step uses $\Delta_{\min}\lesssim\log n$. We complete the proof of (4.3).
Case 2: $\Delta_{\min}\geq C\log n$.
From the decomposition in (D.4), we bound
${1\over n}\left\|\sum_{i=1}^{n}\gamma_{\ell}(X_{i};\omega^{*})(X_{i}-\mu_{\ell}^{*})\right\|_{\Sigma^{*}}\leq~\textrm{I}_{\ell}^{\prime}+\textrm{II}_{\ell}^{\prime}$
where
$\textrm{I}^{\prime}$ $:={1\over n}\left\|\sum_{i\in\widehat{G}_{\ell}}\gamma_{\ell}(X_{i};\omega)(X_{i}-\mu_{\ell}^{*})\right\|_{\Sigma^{*}},\quad\textrm{II}^{\prime}:={1\over n}\left\|\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\gamma_{\ell}(X_{i};\omega)(X_{i}-\mu_{\ell}^{*})\right\|_{\Sigma^{*}}.$
First, note that
$\textrm{I}^{\prime}\leq{1\over n}\left\|\sum_{i\in\widehat{G}_{\ell}}N_{i}\right\|_{2}+{1\over n}\left\|\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\omega^{*}))N_{i}\right\|_{2}$
and Lemma 32 together with $\mathcal{E}_{\pi}$ states that
$\mathbb{P}\left\{{1\over n}\left\|\sum_{i\in\widehat{G}_{\ell}}N_{i}\right\|_{2}\lesssim\sqrt{d\pi_{\ell}^{*}\over n}+\sqrt{\pi_{\ell}^{*}\log n\over n}\right\}\geq 1-n^{-1}.$ (D.13)
Second,
${1\over n}\left\|\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\omega^{*}))N_{i}\right\|_{2}$ $\leq\sup_{v\in\mathbb{S}^{d}}{1\over n}\sqrt{\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\omega^{*}))}\sqrt{\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\omega^{*}))(v^{\top}N_{i})^{2}}$
$\leq{\sqrt{n_{\ell}}\over n}\biggl{\|}\sum_{i\in\widehat{G}_{\ell}}N_{i}N_{i}^{\top}\biggr{\|}_{\rm op}^{1/2}~\max_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\omega^{*})).$
Recalling (C.56) and (C.57). Consider the event
$\mathcal{E}_{\Delta}:=\bigcap_{i=1}^{n}\left\{|N_{i}^{\top}\sigma_{k\ell}(\omega^{*})|\geq{1\over 4}\Delta_{k\ell},~\forall~k\neq\ell\right\}$ (D.14)
which holds with probability at least $1-2nL^{2}\exp(-\Delta_{\min}/32)\geq 1-n^{-1}$ by $\Delta_{\min}\geq C\log n$ and $n\pi_{\min}^{*}\geq C\log(n)(1\vee\log(n)/d)$. On $\mathcal{E}_{\Delta}$, we have that for any $i\in\widehat{G}_{\ell}$,
[…]
Further invoking Lemma 33 and $\mathcal{E}_{\pi}$ gives
${1\over n}\left\|\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\omega^{*}))N_{i}\right\|_{2}\lesssim{\sqrt{n_{\ell}(d+n_{\ell})}\over n\pi_{\ell}^{*}}\exp(-\Delta_{\min}/4)$ (D.16)
with probability at least $1-e^{-cn\pi_{\ell}^{*}}-2n^{-1}\geq 1-3n^{-1}$.
By similar arguments, we also have
$\textrm{II}^{\prime}\leq{1\over n}\sqrt{\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\gamma_{\ell}(X_{i};\omega)}\sup_{v\in\mathbb{S}^{d}}\sqrt{\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\gamma_{\ell}(X_{i};\omega)\langle v,X_{i}-\mu_{\ell}^{*}\rangle_{\Sigma^{*}}^{2}}.$
On $\mathcal{E}_{\Delta}$, for any $i\in\widehat{G}_{a}$ with $a\neq\ell$, recall from (C.67) that
$\gamma_{\ell}(X_{i};\omega^{*})~\leq{\pi_{\ell}^{*}\over\pi_{a}^{*}}\exp(-\Delta_{\min}/4).$ (D.17)
It then follows that on $\mathcal{E}_{\Delta}\cap\mathcal{E}_{\pi}$,
$\textrm{II}^{\prime}$ $\lesssim{1\over n}\sqrt{\sum_{a\in[L]\setminus\{\ell\}}{n_{a}\over\pi_{a}^{*}}}~\sup_{v\in\mathbb{S}^{d}}\sqrt{\sum_{a\in[L]\setminus\{\ell\}}{1\over\pi_{a}^{*}}\sum_{i\in\widehat{G}_{a}}\left[(v^{\top}N_{i})^{2}+\Delta_{a\ell}\right]}~\pi_{\ell}^{*}\exp(-\Delta_{\min}/4)$
$\leq\sqrt{L}~\sup_{v\in\mathbb{S}^{d}}\sqrt{\sum_{a\in[L]\setminus\{\ell\}}{1\over n\pi_{a}^{*}}\sum_{i\in\widehat{G}_{a}}(v^{\top}N_{i})^{2}+L\Delta_{\max}}~\pi_{\ell}^{*}\exp(-\Delta_{\min}/4)$
$\lesssim\sqrt{L}\left(\biggl{\|}{1\over n\pi_{\min}^{*}}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}N_{i}N_{i}^{\top}\biggr{\|}_{\rm op}^{1/2}+\sqrt{L\Delta_{\max}}\right)~\pi_{\ell}^{*}\exp(-\Delta_{\min}/4).$
Invoking Lemma 33 and using $n\geq d$ gives
$\textrm{II}^{\prime}\lesssim\left(\sqrt{L\over\pi_{\min}^{*}}+L\sqrt{\Delta_{\max}}\right)~\pi_{\ell}^{*}\exp(-\Delta_{\min}/4)$ (D.18)
with probability $1-2n^{-1}-e^{-n}$.
Finally, combining Eqs. D.13, D.16 and D.18 and invoking $\mathcal{E}_{\pi}$ concludes that
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\mu_{\ell}^{*}\|_{\Sigma^{*}}$ $={1\over\widehat{\bm{\pi}}_{\ell}(\omega^{*})}{1\over n}\left\|\sum_{i=1}^{n}\gamma_{\ell}(X_{i};\omega^{*})(X_{i}-\mu_{\ell}^{*})\right\|_{\Sigma^{*}}$
$\lesssim\sqrt{d+\log n\over n\pi_{\ell}^{*}}+\left(\sqrt{L\over\pi_{\min}^{*}}+L\sqrt{\Delta_{\max}}+\sqrt{1+{d\over n\pi_{\ell}^{*}}}~{1\over\pi_{\ell}^{*}}\right)\exp\left(-{1\over 4}\Delta_{\min}\right)$
$\lesssim\sqrt{d+\log n\over n\pi_{\ell}^{*}}$ (D.19)
where the last step uses (3.4) and $\Delta_{\min}\geq C\log n$. This proves the second case, hence completes the proof.
@@
and note that
$\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}[H_{i}\mid Y=a]=\Sigma^{*-1/2}\sum_{\ell=1}^{L}\mathbb{E}[\gamma_{\ell}(X;\omega^{*})(X-\mu_{\ell}^{*})(X-\mu_{\ell}^{*})^{\top}]\Sigma^{*-1/2}=\mathbf{I}_{d}.$ (D.21)
As before, we consider two cases.
Case 1: $\Delta_{\min}\leq C\log n$.
By (D.1.3) and (D.21), adding and subtracting terms gives $d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))\leq\textrm{I}+\textrm{II}+\textrm{III}$ where
$\textrm{I}:=$~ ${1\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\left(H_{i}-\mathbb{E}[H_{i}\mid Y_{i}=a]\right)\right\|_{\rm op}$
$\textrm{II}:=$~ $\left\|\sum_{a=1}^{L}\left({n_{a}\over n}-\pi_{a}^{*}\right)\mathbb{E}[H_{i}\mid Y_{i}=a]\right\|_{\rm op}$
$\textrm{III}:=$~ $\left\|\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\omega^{*})\Sigma^{*-1/2}(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))^{\top}\Sigma^{*-1/2}\right\|_{\rm op}.$ (D.22)
In the following we bound each term on the event $\mathcal{E}_{\pi}$ and $\mathcal{E}_{2}$ given by (D.5) and (D.10), respectively.
@@
$\textrm{I}_{1}$ $:={1\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\left(H_{i}1\{\mathcal{E}_{2,i}\}-\mathbb{E}[H_{i}1\{\mathcal{E}_{2,i}\}\mid Y_{i}=a]\right)\right\|_{\rm op},$
$\textrm{I}_{2}$ $:={1\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}\left[H_{i}1\{\mathcal{E}_{2,i}^{c}\}\mid Y_{i}=a\right]\right\|_{\rm op}.$
First, note that
$\textrm{I}_{2}$ $\leq\max_{i\in[n]}\sum_{\ell=1}^{L}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}1\{\mathcal{E}_{2,i}^{c}\}\right]$
$\leq\max_{i\in[n]}\left(2\Delta_{\max}\mathbb{P}(\mathcal{E}_{2,i}^{c})+2\sqrt{\mathbb{E}[\|N_{i}\|_{2}^{4}]}\sqrt{\mathbb{P}(\mathcal{E}_{2,i}^{c})}\right)$
$\lesssim{d+\Delta_{\max}\over n}.$ (D.23)
We then bound $\textrm{I}_{1}$ by the matrix-valued Bernstein inequality of Lemma 35. First, for any $i\in[n]$, almost surely,
$\|H_{i}1\{\mathcal{E}_{2,i}\}\|_{\rm op}$ $\leq\sum_{\ell=1}^{L}\gamma_{\ell}(X_{i};\omega^{*})\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}1\{\mathcal{E}_{2,i}\}$
$\leq 2\|N_{i}\|_{2}^{2}1\{\mathcal{E}_{2,i}\}+2\Delta_{\max}$
$\lesssim d+\log n+\Delta_{\max}.$
@@
$\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}[H_{i}^{2}1\{\mathcal{E}_{2,i}\}]\right\|_{\rm op}$
$\leq\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}\left[\gamma_{a}^{2}(X_{i};\omega^{*})\Sigma^{*-1/2}(X_{i}-\mu_{a}^{*})(X_{i}-\mu_{a}^{*})^{\top}\Sigma^{*-1/2}\|X_{i}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}\mid Y_{i}=a\right]\right\|_{\rm op}$
$\quad+\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}\left[\left(\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\Sigma^{*-1/2}(X_{i}-\mu_{\ell}^{*})(X_{i}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}\right)^{2}\mid Y_{i}=a\right]\right\|_{\rm op}.$
@@
$\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}\left[\gamma_{a}(X_{i};\omega^{*})N_{i}N_{i}^{\top}\|N_{i}\|_{2}^{2}\mid Y_{i}=a\right]\right\|_{\rm op}\leq\sup_{v\in\mathbb{S}^{d}}n\mathbb{E}[(N_{i}^{\top}v)^{2}\|N_{i}\|_{2}^{2}]\leq 3nd.$ (D.24)

### v2

@@ D.1 Proof of Theorem 5: concentration inequalities of the M-steps at the true parameter
Taking the union bounds over $\ell\in[L]$, choosing $t=\log(n)$ and using $n\pi_{\min}^{*}\geq C\log n$ yield (4.2). ∎
@@
Proof.
We work on the event
@@
Pick any $\ell\in[L]$. Notice that
By $\mathbb{E}[V_{\ell,i}]=\mathbb{E}[\gamma_{\ell}(X_{i};\omega^{*})](\bm{\mu}_{\ell}(\omega^{*})-\mu_{\ell}^{*})=0$, we have
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}={1\over n\widehat{\bm{\pi}}_{\ell}(\omega^{*})}\left\|\sum_{i=1}^{n}\left(V_{\ell,i}-\mathbb{E}[V_{\ell,i}]\right)\right\|_{\Sigma^{*}}.$
to which we use a truncation argument together with the vector-valued Bernstein inequality to analyze. From segol2021improved we know $\Sigma^{*-1/2}V_{\ell,i}$ is a sub-Gaussian random vector with some absolute sub-Gaussian constant $C_{v}>0$. Consider the event
$\mathcal{E}_{2}=\bigcap_{i=1}^{n}\mathcal{E}_{2,i},\qquad\mathcal{E}_{2,i}:=\left\{\|V_{\ell,i}\|_{\Sigma^{*}}\leq C_{v}(\sqrt{d}+2\sqrt{\log n})\right\}$ (D.5)
and by Lemma 30, one has $\mathbb{P}(\mathcal{E}_{2,i})\geq 1-n^{-2}$ and $\mathbb{P}(\mathcal{E}_{2})\geq 1-n^{-1}$. We first find that
$\mathbb{E}[\|V_{\ell,i}\|_{\Sigma^{*}}^{2}]$ $\leq\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}\right]$
$\leq\pi_{\ell}^{*}\mathbb{E}\left[\|N_{i}\|_{2}^{2}\right]+2\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}^{*}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})(\|N_{i}\|_{2}^{2}+\Delta_{a\ell})\mid Y_{i}=a\right]$
$\lesssim d\pi_{\ell}^{*}+\sum_{a\in[L]\setminus\{\ell\}}(\pi_{\ell}^{*}+\pi_{a}^{*})(d+\Delta_{a\ell})\exp(-c\Delta_{a\ell})$ by Lemma 21
$\lesssim d\pi_{\ell}^{*}+(L\pi_{\ell}^{*}+1)d\exp{(-c^{\prime}\Delta_{\min})}$ $\text{by }\mathcal{E}_{\pi}$
@@
Then
${1\over n}\left\|\sum_{i=1}^{n}\mathbb{E}[V_{\ell,i}1\{\mathcal{E}_{2,i}^{c}\}]\right\|_{\Sigma^{*}}\leq~\max_{i\in[n]}\mathbb{E}\left[\|V_{\ell,i}\|_{\Sigma^{*}}^{2}\right]^{1/2}\mathbb{P}(\mathcal{E}_{2,i}^{c})^{1/2}\lesssim\sqrt{d\pi_{\ell}^{*}\over n^{2}}.$ (D.6)
@@
${1\over n}\left\|\sum_{i=1}^{n}\Bigl\{V_{\ell,i}1\{\mathcal{E}_{2,i}\}-\mathbb{E}\left[V_{\ell,i}1\{\mathcal{E}_{2,i}\}\right]\Bigr\}\right\|_{\Sigma^{*}}.$ (D.7)
@@
$\max_{i\in[n]}\|V_{\ell,i}1\{\mathcal{E}_{2,i}\}\|_{\Sigma^{*}}\leq C_{v}(\sqrt{d}+2\sqrt{\log n}).$
In conjunction with the above variance bound, applying the vector-valued Bernstein’s inequality in Lemma 33 gives that for all $t\geq 0$,
@@
with probability at least $1-e^{-t}-2n^{-1}$. Together with (D.6), by choosing $t=\log n$ and using $n\pi_{\min}^{*}\geq C\log(n)(1\vee\log(n)/d)$, we obtain that for any $t\geq 1$, with probability $1-3n^{-1}$,
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}$ $\lesssim\sqrt{d\log n\over n\pi_{\ell}^{*}}+{\log n\over n\pi_{\ell}^{*}}\sqrt{d+\log n}\lesssim\sqrt{d\log n\over n\pi_{\ell}^{*}}.$ (D.8)
This completes the proof of (4.3). ∎
@@
and note that $\mathbb{E}[H_{i}]=\mathbf{I}_{d}.$ Adding and subtracting terms gives $d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))\leq\textrm{I}+\textrm{II}$ where
$\textrm{I}:=$~ ${1\over n}\left\|\sum_{i=1}^{n}\left(H_{i}-\mathbb{E}[H_{i}]\right)\right\|_{\rm op}$
$\textrm{II}:=$~ $\left\|\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\omega^{*})\Sigma^{*-1/2}(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))^{\top}\Sigma^{*-1/2}\right\|_{\rm op}.$ (D.10)
By (D.8) and $\mathcal{E}_{\pi}$ in (D.2), with probability at least $1-3n^{-1}$,
II $\lesssim\sum_{\ell=1}^{L}\pi_{\ell}^{*}\left\|\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}^{2}\lesssim{dL\log n\over n}.$ (D.11)
To bound I, note that $\gamma_{\ell}(X_{i};\omega^{*})]^{1/2}(X_{i}-\mu_{\ell}^{*})$ is also sub-Gaussian with some constant $C>0$. In the following we work on the event
$\mathcal{E}_{2}=\bigcap_{i=1}^{n}\mathcal{E}_{2,i},\qquad\mathcal{E}_{2,i}:=\left\{\|[\gamma_{\ell}(X_{i};\omega^{*})]^{1/2}(X_{i}-\mu_{\ell}^{*})\|_{\Sigma^{*}}\leq C_{v}(\sqrt{d}+2\sqrt{\log n})\right\}$
@@
$\textrm{I}_{1}$ $:={1\over n}\left\|\sum_{i=1}^{n}\left(H_{i}1\{\mathcal{E}_{2,i}\}-\mathbb{E}[H_{i}1\{\mathcal{E}_{2,i}\}]\right)\right\|_{\rm op},$
$\textrm{I}_{2}$ $:={1\over n}\left\|\sum_{i=1}^{n}\mathbb{E}\left[H_{i}1\{\mathcal{E}_{2,i}^{c}\}\right]\right\|_{\rm op}.$
We first bound $\textrm{I}_{1}$ by the matrix-valued Bernstein inequality of Lemma 34. For any $i\in[n]$, we have that almost surely,
$\|H_{i}1\{\mathcal{E}_{2,i}\}\|_{\rm op}$ $\leq\sum_{\ell=1}^{L}\gamma_{\ell}(X_{i};\omega^{*})\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}1\{\mathcal{E}_{2,i}\}\lesssim L(d+\log n).$
@@
$\left\|\mathbb{E}[H_{i}^{2}1\{\mathcal{E}_{2,i}\}]\right\|_{\rm op}$
$\leq\left\|\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\gamma_{a}^{2}(X_{i};\omega^{*})\Sigma^{*-1/2}(X_{i}-\mu_{a}^{*})(X_{i}-\mu_{a}^{*})^{\top}\Sigma^{*-1/2}\|X_{i}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}\mid Y_{i}=a\right]\right\|_{\rm op}$
$\quad+\left\|\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\left(\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\Sigma^{*-1/2}(X_{i}-\mu_{\ell}^{*})(X_{i}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}\right)^{2}\mid Y_{i}=a\right]\right\|_{\rm op}.$
@@
$\left\|\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\gamma_{a}(X_{i};\omega^{*})N_{i}N_{i}^{\top}\|N_{i}\|_{2}^{2}\mid Y_{i}=a\right]\right\|_{\rm op}\leq\sup_{v\in\mathbb{S}^{d}}\mathbb{E}[(N_{i}^{\top}v)^{2}\|N_{i}\|_{2}^{2}]\leq 3d.$ (D.12)